Naziah:

Richard’s checking account balance was $57.34 at the beginning of the week. During the week, he recorded the transactions below. A 2-column table with 4 rows. Column 1 is labeled Transaction with entries groceries, credit card bill, deposit, gas. Column 2 is labeled Amount with entries negative 25.37, negative 50, 163.75, negative 13.27. What was Richard’s account balance at the end of the week?

supie:

You have to do 25.37 + 50 + 13.27 to get the expenses of the week (you would have to do Credit card bill+Gas+Groceries to get the expenses of the week [you can add them in any order]) \(25.37 + 50 + 13.27=88.64\)

supie:

So then, we have to find Richard's account balance, we would have to do is \(\text {Richard's account balance at the beginning of the week + Deposit -} \) \(\text{Expenses of the week = Richard's account balance}\), so we plug in the numbers, \(57.34 + 163.75 - 88.64\) and \(57.34 + 163.75 - 88.64=\) \(221.09−88.6=\) \(132.49\)